# High Resolution UDF Meshing via Iterative Networks

### v2

@@ A.7 DoubleCoverUDF [Hou23a] tuning and min-cut
Figure A.8: DoubleCoverUDF min-cut. Comparing reconstructions of DCUDF [Hou23a] with and without the min-cut step at resolution 128 (middle) and 512 (bottom).
@@ A.8 Meshing ground-truth UDF
Although our method is designed to handle imperfect UDFs, we also evaluate it on true UDFs computed directly from ground-truth meshes to verify that it does not introduce unwanted artifacts. We report results on the ShapeNet [Chang15] cars dataset in Tab. A.8 (top), and visualize several triangulations at a resolution of 512 in Fig. A.9. As observed, our method does not introduce significant artifacts; however, as expected, multiple iterations provide no benefit in this setting.
In the same table, we also compute the number of holes as surface boundaries, similarly to the description provided by DCUDF [Hou23a], at resolution 512. We report the average. We notice that the original meshes have a large number of boundaries because they contain multiple detailed components and inner structures. None of the methods faithfully recovers the correct mesh topology. Methods that rely directly on MC triangulation, such as CAP-UDF [Zhou24], NSD-UDF+MC [Stella24] and Ours+MC, tend to suffer from micro-holes and gaps between some of the faces. MeshUDF uses a heuristic specifically designed to reduce this behavior and connect as many portions of the surface as possible, explaining the lower number of holes. DCUDF-T [Hou23a] starts from an inflated mesh, so it generally tends to have fewer holes. Simple postprocessing steps can be applied to all methods to improve the final mesh quality. We take as an example the first of the ShapeNet Cars (object 100715345ee54d7ae38b52b4ee9d36a3), and we apply Trimesh-based postprocessing (fill small holes, merge close vertices, remove spurious faces), showing that all methods improve.
@@
Figure A.9: Triangulating ground-truth UDF. Surface meshing results on the ground-truth UDF in comparison to NSD-UDF [Stella24]. Our method does not introduce significant unwanted artifacts.
Table A.8: Triangulating ground-truth UDFs. Results are reported on the ShapeNet [Chang15] cars dataset, using the true UDF computed from the mesh. We compare against NSD-UDF + MC [Stella24] at multiple resolution with the Median L2 Chamfer Distance $\times 10^{-5}$ with 2M sample points (Chamfer-Distance), and against all MC-based methods with the average number of holes, as surface boundaries at resolution 512(Number of holes). For the latter, we also report results on the first car of the dataset, before and after post-processing the mesh.
@@
ShapeNet cars [Chang15] Car 1 (before post-process.) Car 1 (post-processed)
@@
CAP-UDF [Zhou24] 10024 13265 2125
MeshUDF [Guillard22b] 428 997 240
DCUDF-T [Hou23a] 7 6 6
DCUDF-T-nocut [Hou23a] 2.65 2 2
NSD-UDF+MC [Stella24] 8280 8885 1387

### v3

@@ A.7 DoubleCoverUDF [14] tuning and min-cut
Figure A.8: DoubleCoverUDF min-cut. Comparing reconstructions of DCUDF [14] with and without the min-cut step at resolution 128 (middle) and 512 (bottom).
@@ A.8 Meshing ground-truth UDF
Although our method is designed to handle imperfect UDFs, we also evaluate it on true UDFs computed directly from ground-truth meshes to verify that it does not introduce unwanted artifacts. We report results on the ShapeNet [3] cars dataset in Tab. A.8 (top), and visualize several triangulations at a resolution of 512 in Fig. A.9. As observed, our method does not introduce significant artifacts; however, as expected, multiple iterations provide no benefit in this setting.
In the same table, we also compute the number of holes as surface boundaries, similarly to the description provided by DCUDF [14], at resolution 512. We report the average. We notice that the original meshes have a large number of boundaries because they contain multiple detailed components and inner structures. None of the methods faithfully recovers the correct mesh topology. Methods that rely directly on MC triangulation, such as CAP-UDF [46], NSD-UDF+MC [36] and Ours+MC, tend to suffer from micro-holes and gaps between some of the faces. MeshUDF uses a heuristic specifically designed to reduce this behavior and connect as many portions of the surface as possible, explaining the lower number of holes. DCUDF-T [14] starts from an inflated mesh, so it generally tends to have fewer holes. Simple postprocessing steps can be applied to all methods to improve the final mesh quality. We take as an example the first of the ShapeNet Cars (object 100715345ee54d7ae38b52b4ee9d36a3), and we apply Trimesh-based postprocessing (fill small holes, merge close vertices, remove spurious faces), showing that all methods improve.
@@
Figure A.9: Triangulating ground-truth UDF. Surface meshing results on the ground-truth UDF in comparison to NSD-UDF [36]. Our method does not introduce significant unwanted artifacts.
Table A.8: Triangulating ground-truth UDFs. Results are reported on the ShapeNet [3] cars dataset, using the true UDF computed from the mesh. We compare against NSD-UDF + MC [36] at multiple resolution with the Median L2 Chamfer Distance $\times 10^{-5}$ with 2M sample points (Chamfer-Distance), and against all MC-based methods with the average number of holes, as surface boundaries at resolution 512(Number of holes). For the latter, we also report results on the first car of the dataset, before and after post-processing the mesh.
@@
ShapeNet cars [3] Car 1 (before post-process.) Car 1 (post-processed)
@@
CAP-UDF [46] 10024 13265 2125
MeshUDF [13] 428 997 240
DCUDF-T [14] 7 6 6
DCUDF-T-nocut [14] 2.65 2 2
NSD-UDF+MC [36] 8280 8885 1387